A circular coin has a radius of 3/8 inch.When lying flat, how much area does the coin cover, in square inches?
A.3pi/4 B.3pi/16 C.6pi/16 D.9pi/8

Answer by Alan3354(69443)

A circular coin has a radius of 3/8 inch.When lying flat, how much area does the coin cover, in square inches?
------------
Area = pi*r^2 = pi(3/8)^2
= 9pi/64 sq inches